OpenStudy (anonymous):

Given the function g(x) = -3x + 4, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these two values. The value of g(-2) is smaller than the value of g(4). The value of g(-2) is the same as the value of g(4). The values of g(-2) and g(4) cannot be compared. The value of g(-2) is larger than the value of g(4)

OpenStudy (anonymous):

PLEASE HELP!!!

OpenStudy (solomonzelman):

to compare and contrast g(-2) and g(4) you would first have to find each one of them

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle g(x) = -3x + 4 }\) if I were to find g(3) I would plug in 3 instead of x, into the g(x). \(\large\color{black}{ \displaystyle g(\color{blue}{3}) = -3\color{blue}{(3)}+ 4 }\) \(\large\color{black}{ \displaystyle g(\color{blue}{3}) =-9+ 4 }\) \(\large\color{black}{ \displaystyle g(\color{blue}{3}) =-5 }\)

OpenStudy (solomonzelman):

same way, find g(-2) and g(4). To find g(-2) plug in -2 for x To find g(4) plug in 4 for x

OpenStudy (anonymous):

okay let me do that

OpenStudy (solomonzelman):

sure, go ahead.

OpenStudy (anonymous):

so... g(-2) = 10 and g(4) = -8?

OpenStudy (solomonzelman):

yes, very good. Both are correct!

OpenStudy (solomonzelman):

Now knowing the value of g(-2) and g(4) choose the appropriate answer choice.

OpenStudy (anonymous):

So its the last choice.

OpenStudy (solomonzelman):

Yes, brilliant !

OpenStudy (anonymous):

omg thanks so much for your help!!! i wish i had medals :(

OpenStudy (solomonzelman):

Clever! ty, and yw!

OpenStudy (anonymous):

wait i do have

OpenStudy (anonymous):

i gave you my medal xD

OpenStudy (solomonzelman):

:)

OpenStudy (solomonzelman):

need more help?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Sorry i am bad at math, do I have to post in a different question?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

I will have to depart if not to leave at all though.. sorry for leaving/delay.

OpenStudy (anonymous):

ok